3) find the integral surface of the equation (y-z){2xyp+(x^2-y^2)q}+z(x^2-y^2)=0 through the curve x=t^2, y=0 , z=t^3


1
Expert's answer
2021-05-31T13:15:44-0400

"Question:\\ find \space the\space integral\space surface \space of \space the\space equation \\(y-z){2xyp+(x^2-y^2)q}+z(x^2-y^2)=0\\
 through\space the\space curve\\ x=t^2, y=0 , z=t^3\\
solution:
\\
use\space lagrange's \space method
\\
\frac{dx}{(y-z)2xy}=\frac{dy}{(y-z)(x^2-y^2)}=\frac{dz}{y^2-x^2}\\
Now \space I=II\\
\frac{dx}{(y-z)2xy}=\frac{dy}{(y-z)(x^2-y^2)}\\
{(x^2-y^2)}{dx}-{2xy}{dy}=0\\
this \space is\space exact\space equation\space\\
(since\space if\space Mdx+Ndy=0,\space and\space \frac{∂M}{∂y}=\frac{∂N}{∂x}\space then\space it \space is \space called\space exact \space equation.
Now\space we \space know\space that\\
for\space exact \space equations\space ,solution \space is\space\\
\int Mdx+\int [N-\frac{∂\int Mdx}{∂y}] dy=c_1\\
hence\\
\frac{x^3}{3}-y^2x=c_1.................................(1)\\
Now \space II=III\\
dy=(z-y)dz\\
\frac{dy}{dz}+y=z\\
it \space is \space linear \space diff.\space eq.//
for \space it's \space solution\space \\
I.F.=e^{\int P dz}, \space where \space p=1\space multiple\space of \space y\\
I.F.=e^{\int 1 dz}={e}^{z}\\
and \\
solution\space as\\
{e}^{z}.y=\int {e}^{z}.z \space dz
{e}^{z} .y= {e}^{z}(z-1)+c_2..............................(2)\\

now\space we \space have\space 2\space equations\\
\frac{x^3}{3}-y^2x=c_1.................................(1)\\
{e}^{z} .y= {e}^{z}(z-1)+c_2..............................(2)\\
given \space curve
x=t^2, y=0 , z=t^3;
put \space this\space values\space in\space equations;
\frac{(t^2)^3}{3}-(0)^2(t^2)=c_1\\
c_1=\frac{(t^2)^3}{3}................................(4)\\
and\\

{e}^{z} .y= {e}^{z}(z-1)+c_2\\
{e}^{t^3} .(0)= {e}^{t^3}(t^3-1)+c_2\\
c_2={e}^{t^3}(1-t^3)...........................(5)\\
now\space put\space value\space of \space t^3=\sqrt{3c_1}\space in\space (5)\\
c_2={e}^{\sqrt{3c_1}}(1-\sqrt{3c_1})\\
now\space put\space value\space of \space c_1\space and\space c_2\\
{e}^{z} .y- {e}^{z}(z-1)={e}^{\sqrt{3(\frac{x^3}{3}-y^2x)}}(1-\sqrt{3(\frac{x^3}{3}-y^2x)})\\"